How many moles of nitrogen, N, are in 64.0 g of nitrous oxide, N2O?
1
Expert's answer
2020-05-06T14:00:37-0400

n(N)=(m(N2O)/Mr(N2O))*2=(64/44)*2= 2.91 mol

Answer: 2.91 mol of N